OpenStudy (anonymous):

MEDAL & FAN ANYBODY? Use the function x2 + 2x – 15 = 0 The zeroes of the function are: a. 0 and 0 b. 3 and 5 c. 3 and –5 d. Undefined

OpenStudy (cwrw238):

you can plug in the given values of x and y into the equation and see which ones fit

OpenStudy (anonymous):

d

OpenStudy (cwrw238):

no

OpenStudy (cwrw238):

its not a because 0^2 + 2(0) - 15 = -15

OpenStudy (cwrw238):

its either b or c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i think

OpenStudy (cwrw238):

3^ 2 + 2(3) - 15 =9 + 6 - 15 = 0 now try substituting -5

OpenStudy (cwrw238):

(-5)^2 + 2(-5) - 15 = ?